1/1

Lord George Gordon Byron ,  Lord George Gordon Byron ,
She Walks in Beauty  Sie geht in Schönheit

tradukita de Adolf Böttger

She walks in Beauty, like the Night Of cloudless climes and starry skies; And all that’s best of dark and bright Meet in her aspect and her eyes; Thus mellowed to the tender light Which Heaven to gaudy day denies – Sie geht in Schönheit, gleich der Nacht In wolkenlosem Sternenlicht; Des Schattens und des Lichtes Pracht Eint sich in ihrem Angesicht; Aus dem ein milder Schimmer lacht, Der stets dem grellen Tag gebricht.

One shade the more – one ray the less Had half impaired the nameless grace Which waves in every raven tress Or softly lightens o’er her face Where thoughts serenely sweet express How pure – how dear their dwelling place!

Ein Strahl hinweg, ein Schatten mehr, Und fort würd’ auch die Anmut sein, Die aus dem Rabenlockenmeer Die Stirn umglänzt mit sanftem Schein, Wo die Gedanken süß und hehr Verkünden, dass ihr Wohnsitz rein.

And on that cheek, and o’er that brow, So soft – so calm – yet eloquent, The smiles that win, the tints that glow, But tell of days in goodness spent – A mind at peace with all below – A Heart – whose love is innocent!

Und auf der Stirn, dem Wangenpaar, Spricht von dem reinsten Jugendmut So sanft beredt, so ruhig klar Des Lächelns Reiz, der Farben Glut, Von einem Herzen wunderbar, Wo Liebe voller Unschuld ruht.

Verkinto de tiu ĉi Angla poemo estas Lord George Gordon Byron (∗1788-01-22 – †1824-04-18).

Arg-1109-2238 (2015-01-17 17:34:42)

Tiu ĉi poemo troviĝas en http: // petercochran. files. wordpress. com/ 2009/ 03/ hebrew_ melodies. pdf .

Traduko de la Angla poemo “She Walks in Beauty” de Lord George Gordon Byron (∗1788-01-22 – †1824-04-18) en la Germanan de Adolf Böttger (∗1815-05-21 – †1870-11-16).

Arg-1109-2239 (2015-01-17 17:36:39)

Tiun ĉi poem-germanigon mi, Manfred Retzlaff, trovis en la malnova libro “Byron’s sämmtliche Werke”, presita en la germana urbo Lepsiko (Leipzig) en la jaro 1841. Pri la tradukinto vidu la retejon http: // de. wikipedia. org/ wiki/ Adolf_ BГ�ttger .

http://www.poezio.net/version?poem-id=1109&version-id=,2238,2239